OpenStudy (anonymous):

Please help! Find the third term of the sequence a_1=12, a_(n+1)=(a_n)/(n+1

OpenStudy (apoorvk):

Is this an AP? That thing you wrote for the nth term seems odd - could you check that out please?

OpenStudy (anonymous):

Hmm... Yes, it's the problem word for word. :/

OpenStudy (anonymous):

Can you help or does it not make sense? :/

OpenStudy (apoorvk):

well, \[a_n = \frac{a_{n-1}}{n}\]and \[a_1 =12 ~~\rightarrow given\] so, \[a_2 = a_1/2 = 12 /2\] and hence, \[a_3 = a_2/3\] helps?

OpenStudy (anonymous):

6. THANKS APOORVK!

OpenStudy (apoorvk):

actually it's 12/2/3 = 6/3 = 2. @mathsux4real